OpenStudy (anonymous):

Once again PLEASE HELP<3 Give medals.love.etc Solve 3x – 5y = 15 by the substitution method. y = 2x + 4

OpenStudy (anonymous):

medals or love?

OpenStudy (bahrom7893):

u should be able to do this yourself now

OpenStudy (bahrom7893):

Don't help guys.. let's see if she can do this

OpenStudy (anonymous):

solve 3x – 5y = 15 by the substitution method. y = 2x + 4 ---- Substitute for "y" and solve for "x": 3x-5(2x+4) = 15 3x-10x-20 = 15 -7x = 35 x = -5

OpenStudy (bahrom7893):

aww come on!

OpenStudy (lgbasallote):

GO ISMELL!!!!

OpenStudy (anonymous):

WOWOWOWOW u did that fast

OpenStudy (bahrom7893):

we solved like 5 of the same questions with different numbers..

OpenStudy (lgbasallote):

I mean ILIKEGOODSMELLL =))))

OpenStudy (anonymous):

No, id like help. lol i only have 5 more q's. And @ fool, either or :p

OpenStudy (anonymous):

comon bahrom she wants to marry you ... :P

OpenStudy (lgbasallote):

im really sorry for the offensive mistake of your name =)))))

OpenStudy (anonymous):

@ilikegoodsmells I ANSWER IN TOP!

OpenStudy (anonymous):

Thanks hershey kisses. You are a life saver. Fan time

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

lol ughh i want fan time to :(

OpenStudy (anonymous):

I already have a gf so I don't need the second :P

OpenStudy (bahrom7893):

igba i love my name

OpenStudy (anonymous):

awww thanks hershy :)

OpenStudy (anonymous):

yw :D

OpenStudy (lgbasallote):

it's a small L not a big i @bahrom :) and honetly...your name could be mistaken at first sight =))

OpenStudy (anonymous):

AWW WO FANS YESSS IM IN DISSS

OpenStudy (anonymous):

I love you guys. So much kinder than Oklahoma kids.

OpenStudy (bahrom7893):

really? lol lgba...

OpenStudy (anonymous):

i undrstnd dat u ppl r really eager for V-DAY

OpenStudy (anonymous):

thans igb i swear it was bathroom

OpenStudy (anonymous):

@ilikegoodsmell your other question the answer will be ;) solve 3x – 5y = 15 by the substitution method. y = 2x + 4 ---- Substitute for "y" and solve for "x": 3x-5(2x+4) = 15 3x-10x-20 = 15 -7x = 35 x = -5

OpenStudy (anonymous):

Lol

OpenStudy (lgbasallote):

@tyler...again it's a small L not a big I...it's lgbasallote as in Los Angeles